OpenStudy (anonymous):

Double integral problem. For some reason I just can't get this one. Will post in the first comment to make it neat. Please wait 1 minute.

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\int\limits_{0}^{(6-3x)/2} 6 - 3x - 2y, dydx\]

OpenStudy (anonymous):

I need all the steps so i can see where I went wrong. Thanks.

OpenStudy (anonymous):

Is the second limit of integration (with respect to y) \[\frac{6-3x}{2}?\] The text is a bit small.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Focus on the inner integral first: \[\large\int_{0}^{(6-3x)/2}(6-3x-2y)\;dy\\ \large\left[6y-3xy-y^2\right]_{0}^{(6-3x)/2}\\ \left[6\left(\frac{6-3x}{2}\right)-3x\left(\frac{6-3x}{2}\right)-\left(\frac{6-3x}{2}\right)^2\right]-\left[6(0)-3x(0)-(0)^2\right]\\ 9-9x+\frac{9}{4}x^2\] Now the integral is \[\int_{0}^{2}\left(9-9x+\frac{9}{4}x^2\right)\;dx\] The integral is much simpler here, so I'll leave the rest to you.